# Does GPT Really Get It? A Hierarchical Scale to Quantify Human and AI’s Understanding of Algorithms

### v1

@@ Appendix B Further details on LLM experiments
‘You will answer a series of questions related to the Euclidean algorithm for computing the greatest common denominator (gcd) of two integers. You provide concise responses and do not include detail or explanations unless explicitly requested by the user.’
@@
We conducted two experiments with GPT. In the first, GPT was asked five randomized versions of the questions in Section A (a quiz), with the images being replaced with text descriptions of the graphs. For each quiz, the questions corresponding to each level of understanding were asked in order, and the previous questions and responses were included in the API query. This quiz was repeated thirty times for each version. Three quizzes were discarded due to the responses to the randomized questions being incomparable with the others. All trials were hand-graded on a scale from zero to two using the scale described in Section 3.

### v3

@@ Appendix B Further details on LLM experiments
‘You will answer a series of questions related to the Euclidean algorithm for computing the greatest common divisor (gcd) of two integers. You provide concise responses and do not include detail or explanations unless explicitly requested by the user.’
@@
We conducted two experiments with GPT. In the first, GPT was asked randomized versions of the questions given to students (a quiz), with the images being replaced with text descriptions of the graphs. For each quiz, the questions corresponding to each of the levels of understanding were asked in order, and the previous questions and responses were included in the API query. This quiz was repeated thirty times for each GPT version and algorithm. All trials were hand-graded on a scale from zero to two.
@@
Appendix C Additional Results
Figures 8 and 9 show the mean score per question number (originally presented in Figures 2 and 3) split by algorithm.
Figure 8: The average scores across students who self-reported that they understood the algorithm on the Euclidean algorithm (left) and Ford-Fulkerson algorithm (right). Error bars show the 95% confidence interval.
Figure 9: The average scores between versions of GPT on the Euclidean algorithm (left) and Ford-Fulkerson algorithm (right). Error bars show the 95% confidence interval.
In Table 10, we include the mean performance for each of the seven types of questions in the five studied populations (Graduate, Undergraduate, and the three GPT versions).
GCD GPT 3.5 GPT 4 GPT 4o UG Grad
Q1 2.00 2.00 2.00 2.00 2.00
Q2 1.23 1.57 1.70 1.50 2.00
Q3 1.93 2.00 2.00 1.25 2.00
Q4a 0.40 1.67 1.57 0.50 1.83
Q5a 0.67 1.80 1.77 0.50 2.00
Q4b 1.00 1.93 1.91 0.67 2.00
Q5b 0.17 1.47 1.17 0.75 2.00
FLOW GPT 3.5 GPT 4 GPT 4o UG Grad
Q1 1.00 1.97 2.00 2.00 2.00
Q2 0.43 1.27 1.30 2.00 1.60
Q3 0.80 1.57 2.00 1.00 1.20
Q4a 0.17 0.63 0.57 1.00 1.60
Q5a 0.07 0.13 0.37 0.00 1.40
Q4b 0.76 1.76 1.69 1.33 0.67
Q5b 0.77 1.77 1.73 1.33 1.33
Figure 10: Mean score out of two by population on the two tested algorithms, Euclidean (Top) and Ford Fulkerson (Bottom). The highest score(s) for each question and each algorithm are highlighted.